I'm trying to solve for a point of intersection between two functions. I think that to solve for the x-intercept, you have to set the two equations equal to each other and solve, but I'm not sure how to solve the two functions that I have. They are (x/60)*15+38 and (x/60)*7+46
Any help with this problem would be greatly appreciated!!!

Answer by Alan3354(69443)

(x/60)*15+38 and (x/60)*7+46
They're functions in x.
f(x) = (x/60)*15 + 38 = (x/4) + 38
g(x) = (7x/60) + 46
They intersect where they're equal.
(x/4) + 38 = (7x/60) + 46
x/4 = 7x/60 + 8
15x = 7x + 480
8x = 480
x = 60
------
f(60) = 53
g(60) = 53
----------
Plotted on an x-y coordinate system, they intersect at (60,53)